# On a generalisation of Cameron’s base size conjecture

### v2

@@ 9.2. Unipotent elements
Let $x\in\bar{G}_{\sigma}$ be a unipotent element and let $W$, $W_{\bar{H}}$, and $\Omega$ be as in Section 9.1. Note that we again have $|C_{\Omega}(x)|=\chi(x)$, where $\chi=1^{\bar{G}_{\sigma}}_{\bar{H}_{\sigma}}$. Now if $\widehat{W}$ denotes the set of (ordinary) irreducible characters of $W$, then [51, 2.4] gives
@@
and $R_{\phi}(x)$ are the Foulkes functions of $\bar{G}_{\sigma}$. The values $n_{\phi}$ are recorded in [51, pp. 413–415] when $\bar{G}_{\sigma}$ is untwisted and $\bar{H}$ is a maximal parabolic subgroup of $\bar{G}$, and we can find them in [27, Table 8] if $G_{0}=E_{6}(q)$ and $H=P_{1,6}$, or $P_{3,5}$, or $G_{0}=F_{4}(q)$ and $H=P_{1,4}$ or $P_{2,3}$. Finally, if $G_{0}={}^{2}E_{6}(q)$, then the values $n_{\phi}$ are recorded in [25, p. 14].
The values $R_{\phi}$ of unipotent elements of order $p$ have been computed by Lübeck [64] and are given to us in GAP-readable form, so we can compute ${\rm fpr}(x,G/H)$ precisely for all groups with $G_{0}\in\mathcal{E}_{2}\cup\mathcal{E}_{3}$, all unipotent elements and all maximal parabolic subgroups of $G$. We take the opportunity to thank him for making this data available to us.
@@ 9.3. Field and graph-field automorphisms
Let $x\in G$ be a field automorphism of prime order $r$. We write $\bar{G}_{\sigma}=\bar{G}(q)$, $\bar{H}_{\sigma}=\bar{H}(q)$ and $C_{\bar{G}_{\sigma}}(x)=\bar{G}^{\epsilon}(q^{1/r})$. As explained in the proof of [51, 6.1], we have $x^{\bar{G}_{\sigma}}\cap\bar{H}_{\sigma}x=x^{\bar{H}_{\sigma}}$, and moreover $C_{\bar{H}_{\sigma}}=\bar{H}^{\epsilon}(q^{1/r})$, where $\bar{H}^{\epsilon}(q^{1/r})$ is the corresponding parabolic subgroup of $C_{\bar{G}_{\sigma}}(x)$ and $\epsilon=$- if $x$ is a graph-field automorphism and $\epsilon=+$ otherwise. In particular, by inspecting the proof of [51, 6.1], we get the following statement:
@@ 9.4. Graph automorphisms
Here $G_{0}=E_{6}^{\epsilon}(q)$ and there are two classes of graph automorphisms, namely those of type $F_{4}$, and those of type $C_{4}$. We record an upper bound $t(H,\tau,q)$ for every $H\in\mathcal{P}_{G}$ and every involutory graph automorphism $\tau\in G$. If $p=2$, then [5, Section 19] implies that either $C_{\bar{G}}(x)=F_{4}$ or $C_{F_{4}}(t)$ for a long root element $t\in F_{4}$. The value $t(H,\tau,q)$ can be obtained via [51, 2.6] when $C_{\bar{G}}(x)=F_{4}$, and if $G_{\bar{G}}(x)=C_{F_{4}}(t)$, then a bound $t(H,\tau,q)$ is recorded in [25, Table 6]. For the reader’s convenience, we record these values in Table 17.
@@
If $p>2$, then $C_{\bar{G}}(\tau)=F_{4}$ or $C_{4}$. Sufficient fixed point ratio bounds can be obtained by working as in the proof of [51, 6.4]. In the following proposition we carry out these computations to obtain $t(H,\tau,q)$. Note that we will only need an upper bound over all parabolic subgroups for our purposes, but those bounds might be of independent interest, and so we record separate upper bounds for each choice of parabolic subgroup. The value $t(H,\tau,q)$ for each distinct pair $(H,\tau)$ is recorded in Table 18.
@@
First observe that $H=N_{G}(P)$, where $P$ is a $\tau$-stable parabolic subgroup of $\bar{G}=E_{6}$. Moreover, as shown in [51, 6.4], $C_{P}(\tau)$ is a parabolic subgroup of $C_{\bar{G}}(\tau)$. Write $P=QL$, where $Q$ is the unipotent radical of $P$ and $L$ is a Levi subgroup, and we treat each possibility for $P$ in turn.
@@
Next assume that $P=P_{4}$, in which case $L=A_{2}^{2}A_{1}T_{1}$. In this case, $\tau$ must swaps the two $A_{2}$ factors of $L^{\prime}$ and it either centralises the $A_{1}$ factor, or acts on it as an involution with centraliser $T_{1}$. It follows that $C_{L^{\prime}}(\tau)$ is either $A_{2}A_{1}$ or $A_{2}T_{1}$.
@@
Here $q=3^{2m+1}$ for some non-negative integer $m$. In view of the isomorphism $^{2}G_{2}(3)^{\prime}\cong{\rm L}_{2}(8)$, we may assume that $q\geqslant 27$. In this case we can infer from [51, Theorem 1] that ${\rm fpr}(x,G/H)\leqslant(q^{2}-q+1)^{-1}$ for all prime order elements $x\in G$ and $H\in\mathcal{M}_{G}$, and therefore
@@
Here $q=2^{m+1}$ for some $m\geqslant 1$. Observe that [51, Theorem 1] asserts that ${\rm fpr}(x,G/H)\leqslant(q^{2/3}+1)/(q^{2}+1)$ for all prime order elements $x\in G$ and $H\in\mathcal{M}_{G}$, and thus
@@
As noted in [25, Section 4.6], we can infer from [70] that $G$ has two classes of involutions with respective representatives $t_{2}$ and $t^{\prime}_{2}$, where
@@
Moreover, [51, Theorem 1] implies that ${\rm fpr}(x,G/H)\leqslant q^{-4}$ for all involutions $x\in G$ and all $H\in\mathcal{M}_{G}$, so the contribution to $\widehat{Q}(G,\tau)$ from unipotent elements is less than $2q^{14}\cdot(q^{-4})^{6}=2q^{-10}$.
On the other hand, if $x\in G$ is semisimple or a field automorphism of prime order, then [51, Theorem 2] gives ${\rm fpr}(x,G/H)\leqslant 2q^{-6}$, so we have
@@
Let $x\in G$ be an element of prime order $r$. We claim that if ${\rm dim}\,x^{\bar{G}}>20$ or $x$ is a field automorphism of prime order $r\geqslant 5$, then ${\rm fpr}(x,G/H)<(q^{2}(q^{3}-2))^{-1}$ for any $H\in\mathcal{M}_{G}$. If $H$ is parabolic, then the claim follows directly by [51, Theorem 2], so assume that $H$ is non-parabolic. If $H_{0}={}^{3}D_{4}(q^{1/2})$, then we verify that the bound holds by inspecting the proof of [25, 4.42], and in all other cases $|H_{0}|<q^{14}$, so the bound is again satisfied.
Now assume that ${\rm dim}\,x^{\bar{G}}\leqslant 20$, or $r\in\{2,3\}$ and $x$ is either a field or triality graph automorphism. In this case [51, Theorem 1] implies that ${\rm fpr}(x,G/H)<(q^{4}-q^{2}+1)^{-1}$ and we claim that there are at most $3q^{22}$ elements in $G$ with ${\rm dim}\,x^{\bar{G}}\leqslant 20$.
By inspecting [37, Table 4.4], we deduce that there are at most $q^{2}+q+1$ such semisimple classes, and we find that there are $4$ such unipotent classes by inspecting [71]. Moreover, if $x$ is a field automorphism, then either $r=2$ and we have $|x^{G}|<2q^{14}$, or $r=3$ and $|x^{G}|<2q^{56/3}<q^{20}$. Note that $G$ has at most two classes of order $3$ field automorphisms. Finally there are fewer than $4q^{20}+2q^{14}$ triality graph automorphisms, so in total we get
@@
In view of the isomorphism $G_{2}(2)^{\prime}\cong{\rm U}_{3}(3)$, we may assume that $q\geqslant 3$. In particular, $G_{0}=G_{2}(q)$. Let $x\in G$ be a prime order element. If $x$ is an involutory field or graph field automorphism, or a long root element, then [51, Theorem 1] asserts that ${\rm fpr}(x,G/H)\leqslant(q^{2}-q+1)^{-1}$ for any $H\in\mathcal{M}_{G}$ and we can check that there are at most $q^{8}$ such elements in $G$. For all other elements, we claim that ${\rm fpr}(x,G/H)\leqslant(q^{2}(q-1))^{-1}$.
First suppose that $x$ is a field automorphism of order $r\geqslant 3$. If $|H_{0}|\leqslant q^{6}$, then the claim trivially holds, since $|x^{G}|>\frac{1}{2}q^{28/3}$ as noted in [25, 4.29], and so
@@
Next, if $H$ is of type ${\rm SL}_{3}^{\epsilon}(q)$ or $^{2}G_{2}(q)$, then we confirm the required bound holds by inspecting the proofs of [25, 4.30, 4.31] respectively, and if $H$ is of type $G_{2}(q^{1/2})$, then $x$ acts on $H_{0}$ as a field automorphism, so we obtain
@@
which is sufficient. For all other possibilities for $x$, the claim follows immediately from [51, Theorem 2].
@@
Let $\mathcal{P}=\mathcal{S}\cup\mathcal{U}\cup\mathcal{F}$, where $\mathcal{S}$ denotes the set of semisimple elements of prime order in $G$, $\mathcal{U}$ denotes the class of prime order unipotent elements in $G$, and $\mathcal{F}$ denotes the prime order graph, field, and graph field automorphisms of $G$, then we may write
@@
Our aim is to derive suitable bounds on $\alpha$, $\beta$ and $\gamma$ to prove Theorem 7.1 for $G$, which is the strategy that we will be adopting for all groups with socle in $\mathcal{E}_{2}$.
@@
Let us now record an important remark about semisimple elements. Note that Fleischmann and Janiszczak work with simply connected groups of type $E_{6}$ and $E_{7}$ [39], whereas we are interested in adjoint type groups. However, for a semisimple element $x\in\bar{G}_{\sigma}$, we are only using their results on centraliser orders to compute ${\rm fpr}(x,G/H)$. As described in Section 9, we do this by appealing to Proposition 9.1, and so we are only interested in $C_{\bar{G}_{\sigma}}(x)^{0}$, which is the same both in the adjoint and in the simply connected case.
@@
We first consider semisimple elements. Recall that the possible centralisers of semisimple elements in $\bar{G}_{\sigma}$ are parametrised by pairs $(J,[w])$, where $J$ and $w$ are defined in Section 9.1. The possibilities for $(J,[w])$ are listed in [39], and the number of $\bar{G}_{\sigma}$-classes with centraliser type corresponding to a fixed pair $(J,[w])$ is recorded in [65]. Therefore, if $\{x_{1},\ldots,x_{t}\}$ is a complete set of representatives of $\bar{G}_{\sigma}$-classes of each centraliser type and $c_{i}$ denotes the number of $\bar{G}_{\sigma}$-classes corresponding to a pair $(J,[w])$, then we can write
@@
and check that it is less than $q^{-1}$ for all values of $q\geqslant 3$.
Let us see an example. Assume $\epsilon=+$ and consider the $\bar{G}_{\sigma}$-class corresponding to the pair $(J,[w])$ with $J=\{\alpha_{0},\alpha_{2},\alpha_{5},\alpha_{6}\}$ and $w=a_{1}a_{2}$ as given in [39]. Let $y$ be a representative of a $\bar{G}_{\sigma}$-class corresponding to this pair. Inspecting [39] tells us that $C_{\bar{G}_{\sigma}}(y)$ is of type $A_{2}^{2}T_{2}$, and in particular, $|C_{\bar{G}_{\sigma}}(y)|=|{\rm SL}_{3}(q)|^{2}\Phi_{3}(q)$, where $\Phi_{3}(q)=q^{2}-q+1$ denotes the third cyclotomic polynomial. We now turn to [65] and we find that the entry corresponding to the class of $y$ is the one with label $(i,j,k)=(12,1,5)$, and from there we can read the number of classes.
@@
$\beta\leqslant\sum_{i=1}^{r}|x_{i}^{\bar{G}_{\sigma}}|\cdot F(x_{i},q)^{6}$
where $x_{1},\ldots,x_{r}$ represent the prime order unipotent classes of $G$. In this way, we find that
@@
We finally compute a bound on $\gamma$. Let $\gamma_{1}$ denote the contribution from involutory field and graph-field automorphisms, $\gamma_{2}$ denote the odd order field automorphisms and $\gamma_{3}$ denote the contribution from graph automorphisms.
@@
where we recall that $\bar{H}(q)=H_{\sigma}$ and $\bar{H}(q^{1/r})=C_{\bar{H}_{\sigma}}(x)$. Note that this implies that ${\rm fpr}(x,G/H)$ is maximised when $|\bar{H}(q)|$ (and consequently $|\bar{H}(q^{1/r})|$) is maximised, and we certainly have $|\bar{H}(q)|\leqslant|P_{1}(q)|$, where $P_{1}$ is the $P_{1}$-parabolic subgroup of $E_{6}$. Moreover, if $r=2$, then we note that the fixed point ratio of a graph-field automorphism is greater than the fixed point ratio of a field automorphism. We therefore find
$\gamma_{1}\leqslant\frac{3+\epsilon}{2}\cdot\left(\frac{|E_{6}^{\epsilon}(q)|}{|{}^{2}E_{6}(q^{1/2})|}\right)^{-5}\left(\frac{|P_{1}(q)|}{|P_{1}(q^{1/r})|}\right)^{6}$
@@
Finally, we infer from Propositions 8.1, 8.2, and 9.5 that if $x$ is an involutory graph automorhpism, then $f(x,q)\leqslant q^{-5}$ if $C_{\bar{G}}(x)=F_{4}$ and $f(x,q)\leqslant 12q^{-10}$ if $C_{\bar{G}}(x)=C_{4}$. If $q>4$, then we find
@@
If $q\geqslant 5$, then one checks that
@@
Finally, if $q=3$ or $4$, then by only considering the contribution by involutory graph automorphisms (and field and graph-field automorphisms if $q=4$), we easily compute $\gamma\leqslant 0.025$ using Magma.

### v3

@@ 9.1. Semisimple elements
In the process of extracting results from [40], we spotted some minor errors in some places. We take the opportunity to correct them here in case this is useful to the reader.
Let $G$ denote the simply connected group of type $E_{7}$, and let $x\in G$ be a semisimple element corresponding to the pair $(J,[w])$. First consider the case $J=\{\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{4},\alpha_{6}\}$, so that $C_{G}(x)$ is of type $A_{2}^{2}A_{1}$. If $w=w_{2}=a_{5}$, as given in [40, p.127], then we need to add an extra ${\rm SU}_{3}(q)$ factor in $C_{G}(x)$, so that $C_{G}(x)={\rm SU}_{3}(q)^{2}\times{\rm SL}_{2}(q)\times(q+1)^{2}$.
Next suppose that $J=\{\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{5},\alpha_{6},\alpha_{7}\}$, so that $C_{G}(x)$ is of type $A_{3}^{2}$. In this case there is a mismatch in [40] between the $w_{i}$ and the right centraliser structure for $w=w_{i}$ for $i\in\{4,6,7,8\}$, which can be corrected by an appropriate permutation of the $w_{i}$. We record the correct assignment of centraliser to each $w_{i}$ in Table 18.
[…]
Table 18. Correction to [40] for $J=\{\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{5},\alpha_{6},\alpha_{7}\}$.
Finally, suppose that $C_{G}(x)$ is of type $(A_{1}^{3})_{1}$. In this case, by inspecting the Dynkin diagram of $G$, we note that the choice of $J$ in [40] is not correct. One could set $J=\{\alpha_{0},\alpha_{3},\alpha_{7}\}$. We then turn into Magma and run the following code:
G := GroupOfLieType("E7", 3);
W := WeylGroup(G);
J := {3, 7, 126};
WJ:=ReflectionSubgroup(W, J);
Q, f := quo<Normalizer(W, WJ)| WJ>;
Z := [];
cl := Classes(Q);
for i in [1 .. #cl] do
  Append(~Z, cl[i][3]@@f);
end for;
to obtain a list of coset representatives in $N_{W}(W_{J})$ of the $\sigma$-conjugacy classes of $N_{W}(W_{J})/W_{J}$. The final step is to write those representatives as products of $a_{1},\ldots,a_{62}$, as defined in [40, pp.118-119]. To do this, we construct $a_{1},\ldots,a_{62}$ in Magma as permutations and we store them in a list A. We then turn into GAP and we input the lists of permutations $A$ and $Z$ we have obtained. We then run the following code:
G := Group(A);
C := [];
for i in [1 .. Length(Z)] do
Add(C, Factorization(G, Z[i]));
od;
and we obtain the following list of $w_{i}$:
$w_{1}$ $={\rm id},w_{2}=a_{38}a_{22}a_{12},w_{3}=a_{10}a_{5},w_{4}=a_{61}a_{18},w_{5}=a_{5},$
$w_{6}$ $=a_{15}a_{10}a_{5},w_{7}=a_{15}a_{11},w_{8}=a_{15},w_{9}=a_{21}a_{26}a_{21}a_{6},w_{10}=a_{38}a_{15},$
$w_{11}$ $=a_{49}a_{37}a_{15}a_{5},w_{12}=a_{33}a_{60}a_{16}a_{14},w_{13}=a_{53}a_{37}a_{53},w_{14}=a_{49}a_{37},w_{15}=a_{53}a_{55}a_{61}a_{3},$
$w_{16}$ $=a_{49}a_{37}a_{11},w_{17}=a_{55}a_{30}a_{49},w_{18}=a_{36}a_{37}a_{18},w_{19}=a_{49}a_{36}a_{37}a_{15},w_{20}=a_{16}a_{49}a_{26}a_{6}$
Taking $C_{G}(x)$ as in [40, p.122] for each $i\in\{1,\ldots,20\}$, we obtain a complete list of the possible pairs $(J,[w])$ when $C_{G}(x)$ of type $(A_{1}^{3})_{1}$.
@@ 9.2. Unipotent elements
Let $x\in\bar{G}_{\sigma}$ be a unipotent element and let $W$, $W_{\bar{H}}$, and $\Omega$ be as in Section 9.1. Note that we again have $|C_{\Omega}(x)|=\chi(x)$, where $\chi=1^{\bar{G}_{\sigma}}_{\bar{H}_{\sigma}}$. Now if $\widehat{W}$ denotes the set of (ordinary) irreducible characters of $W$, then [52, 2.4] gives
@@
and $R_{\phi}(x)$ are the Foulkes functions of $\bar{G}_{\sigma}$. The values $n_{\phi}$ are recorded in [52, pp. 413–415] when $\bar{G}_{\sigma}$ is untwisted and $\bar{H}$ is a maximal parabolic subgroup of $\bar{G}$, and we can find them in [28, Table 8] if $G_{0}=E_{6}(q)$ and $H=P_{1,6}$, or $P_{3,5}$, or $G_{0}=F_{4}(q)$ and $H=P_{1,4}$ or $P_{2,3}$. Finally, if $G_{0}={}^{2}E_{6}(q)$, then the values $n_{\phi}$ are recorded in [26, p. 125].
The values $R_{\phi}$ of unipotent elements of order $p$ have been computed by Lübeck [65] and are given to us in GAP-readable form, so we can compute ${\rm fpr}(x,G/H)$ precisely for all groups with $G_{0}\in\mathcal{E}_{2}\cup\mathcal{E}_{3}$, all unipotent elements and all maximal parabolic subgroups of $G$.
@@ 9.3. Field and graph-field automorphisms
Let $x\in G$ be a field automorphism of prime order $r$. We write $\bar{G}_{\sigma}=\bar{G}(q)$, $\bar{H}_{\sigma}=\bar{H}(q)$ and $C_{\bar{G}_{\sigma}}(x)=\bar{G}^{\epsilon}(q^{1/r})$. As explained in the proof of [52, 6.1], we have $x^{\bar{G}_{\sigma}}\cap\bar{H}_{\sigma}x=x^{\bar{H}_{\sigma}}$, and moreover $C_{\bar{H}_{\sigma}}=\bar{H}^{\epsilon}(q^{1/r})$, where $\bar{H}^{\epsilon}(q^{1/r})$ is the corresponding parabolic subgroup of $C_{\bar{G}_{\sigma}}(x)$ and $\epsilon=$- if $x$ is a graph-field automorphism and $\epsilon=+$ otherwise. In particular, by inspecting the proof of [52, 6.1], we get the following statement:
@@ 9.4. Graph automorphisms
Here $G_{0}=E_{6}^{\epsilon}(q)$ and there are two classes of graph automorphisms, namely those of type $F_{4}$, and those of type $C_{4}$. We record an upper bound $t(H,\tau,q)$ for every $H\in\mathcal{P}_{G}$ and every involutory graph automorphism $\tau\in G$. If $p=2$, then [5, Section 19] implies that either $C_{\bar{G}}(x)=F_{4}$ or $C_{F_{4}}(t)$ for a long root element $t\in F_{4}$. The value $t(H,\tau,q)$ can be obtained via [52, 2.6] when $C_{\bar{G}}(x)=F_{4}$, and if $C_{\bar{G}}(x)=C_{F_{4}}(t)$, then a bound $t(H,\tau,q)$ is recorded in [26, Table 7]. For the reader’s convenience, we record these values in Table 19.
@@
If $p>2$, then $C_{\bar{G}}(\tau)=F_{4}$ or $C_{4}$. Sufficient fixed point ratio bounds can be obtained by working as in the proof of [52, 6.4]. In the following proposition we carry out these computations to obtain $t(H,\tau,q)$. Note that we will only need an upper bound over all parabolic subgroups for our purposes, but those bounds might be of independent interest, and so we record separate upper bounds for each choice of parabolic subgroup. The value $t(H,\tau,q)$ for each distinct pair $(H,\tau)$ is recorded in Table 20.
@@
First observe that $H=N_{G}(P)$, where $P$ is a $\tau$-stable parabolic subgroup of $\bar{G}=E_{6}$. Moreover, as shown in [52, 6.4], $C_{P}(\tau)$ is a parabolic subgroup of $C_{\bar{G}}(\tau)$. We write $P=QL$, where $Q$ is the unipotent radical of $P$ and $L$ is a Levi subgroup, and we treat each possibility for $P$ in turn.
@@
Next assume that $P=P_{4}$, in which case $L=A_{2}^{2}A_{1}T_{1}$. Here, $\tau$ swaps the two $A_{2}$ factors of $L^{\prime}$ and it either centralises the $A_{1}$ factor, or acts on it as an involution with centraliser $T_{1}$. It follows that $C_{L^{\prime}}(\tau)$ is either $A_{2}A_{1}$ or $A_{2}T_{1}$.
@@
Here $q=3^{2m+1}$ for some non-negative integer $m$. In view of the isomorphism $^{2}G_{2}(3)^{\prime}\cong{\rm L}_{2}(8)$, we may assume that $q\geqslant 27$. In this case we can infer from [52, Theorem 1] that ${\rm fpr}(x,G/H)\leqslant(q^{2}-q+1)^{-1}$ for all prime order elements $x\in G$ and $H\in\mathcal{M}_{G}$, and therefore
@@
Here $q=2^{m+1}$ for some $m\geqslant 1$. Observe that [52, Theorem 1] asserts that ${\rm fpr}(x,G/H)\leqslant(q^{2/3}+1)/(q^{2}+1)$ for all prime order elements $x\in G$ and $H\in\mathcal{M}_{G}$, and thus
@@
As noted in [26, Section 4.6], we can infer from [73] that $G$ has two classes of involutions with respective representatives $t_{2}$ and $t^{\prime}_{2}$, where
@@
Moreover, [52, Theorem 1] implies that ${\rm fpr}(x,G/H)\leqslant q^{-4}$ for all involutions $x\in G$ and all $H\in\mathcal{M}_{G}$, so the contribution to $\widehat{Q}(G,\tau)$ from unipotent elements is less than $2q^{14}\cdot(q^{-4})^{6}=2q^{-10}$.
On the other hand, if $x\in G$ is semisimple or a field automorphism of prime order, then [52, Theorem 2] gives ${\rm fpr}(x,G/H)\leqslant 2q^{-6}$, so we have
@@
Let $x\in G$ be an element of prime order $r$. We claim that if ${\rm dim}\,x^{\bar{G}}>20$ or $x$ is a field automorphism of prime order $r\geqslant 5$, then ${\rm fpr}(x,G/H)<(q^{2}(q^{3}-2))^{-1}$ for any $H\in\mathcal{M}_{G}$. If $H$ is parabolic, then the claim follows directly by [52, Theorem 2], so assume that $H$ is non-parabolic. If $H_{0}={}^{3}D_{4}(q^{1/2})$, then we verify that the bound holds by inspecting the proof of [26, 4.43], and in all other cases $|H_{0}|<q^{14}$, so the bound is again satisfied.
Now assume that ${\rm dim}\,x^{\bar{G}}\leqslant 20$, or $r\in\{2,3\}$ and $x$ is either a field or triality graph automorphism. In this case [52, Theorem 1] implies that ${\rm fpr}(x,G/H)<(q^{4}-q^{2}+1)^{-1}$ and we claim that there are at most $3q^{22}$ elements in $G$ with ${\rm dim}\,x^{\bar{G}}\leqslant 20$.
By inspecting [38, Table 4.4], we deduce that there are at most $q^{2}+q+1$ such semisimple classes, and we find that there are $4$ such unipotent classes by inspecting [74]. Moreover, if $x$ is a field automorphism, then either $r=2$ and we have $|x^{G}|<2q^{14}$, or $r=3$ and $|x^{G}|<2q^{56/3}<q^{20}$. Note that $G$ has at most two classes of order $3$ field automorphisms. Finally there are fewer than $4q^{20}+2q^{14}$ triality graph automorphisms, so in total we get
@@
In view of the isomorphism $G_{2}(2)^{\prime}\cong{\rm U}_{3}(3)$, we may assume that $q\geqslant 3$. In particular, $G_{0}=G_{2}(q)$. Let $x\in G$ be a prime order element. If $x$ is an involutory field or graph field automorphism, or a long root element, then [52, Theorem 1] asserts that ${\rm fpr}(x,G/H)\leqslant(q^{2}-q+1)^{-1}$ for any $H\in\mathcal{M}_{G}$ and we can check that there are at most $q^{8}$ such elements in $G$. For all other elements, we claim that ${\rm fpr}(x,G/H)\leqslant(q^{2}(q-1))^{-1}$.
First suppose that $x$ is a field automorphism of order $r\geqslant 3$. If $|H_{0}|\leqslant q^{6}$, then the claim trivially holds, since $|x^{G}|>\frac{1}{2}q^{28/3}$ as noted in [26, 4.30], and so
@@
Next, if $H$ is of type ${\rm SL}_{3}^{\epsilon}(q)$ or $^{2}G_{2}(q)$, then we confirm the required bound holds by inspecting the proofs of [26, 4.31, 4.32] respectively, and if $H$ is of type $G_{2}(q^{1/2})$, then $x$ acts on $H_{0}$ as a field automorphism, so we obtain
@@
which is sufficient. For all other possibilities for $x$, the claim follows immediately from [52, Theorem 2].
@@
Let $\mathcal{P}=\mathcal{S}\cup\mathcal{U}\cup\mathcal{F}$, where $\mathcal{S}$ denotes the set of semisimple elements of prime order in $G$, $\mathcal{U}$ denotes the class of prime order unipotent elements in $G$, and $\mathcal{F}$ denotes the prime order graph, field, and graph-field automorphisms of $G$. We may then write
@@
Our aim is to derive suitable bounds on $\alpha$, $\beta$ and $\gamma$ to prove Theorem 7.1 for $G$.
@@
Let us now record an important remark about semisimple elements. Note that Fleischmann and Janiszczak work with simply connected groups of type $E_{6}$ and $E_{7}$ in [40], whereas we are interested in the adjoint type groups. However, for a semisimple element $x\in\bar{G}_{\sigma}$, we are only using their results on centraliser orders to estimate ${\rm fpr}(x,G/H)$. As described in Section 9, we do this by appealing to Proposition 9.1, and so we are only interested in $(C_{\bar{G}}(x)^{0})_{\sigma}$, which is the same both in the adjoint and in the simply connected case.
@@
We first consider semisimple elements. Recall that the possible centralisers of semisimple elements in $\bar{G}_{\sigma}$ are parametrised by pairs $(J,[w])$, where $J$ and $w$ are defined in Section 9.1. The possibilities for $(J,[w])$ are listed in [40], and the number of $\bar{G}_{\sigma}$-classes with centraliser type corresponding to a fixed pair $(J,[w])$ is recorded in [66]. Therefore, if $\{x_{1},\ldots,x_{t}\}$ is a complete set of representatives of $\bar{G}_{\sigma}$-classes of each centraliser type corresponding to a fixed pair $(J,[w])$ and $c_{i}$ denotes the number of $\bar{G}_{\sigma}$-classes corresponding to a pair $(J,[w])$, then we can write
@@
and check that it is less than $q^{-1}$ for all $q\geqslant 3$.
Let us see an example. Assume $\epsilon=+$ and consider the $\bar{G}_{\sigma}$-class corresponding to the pair $(J,[w])$ with $J=\{\alpha_{0},\alpha_{2},\alpha_{5},\alpha_{6}\}$ and $w=a_{1}a_{2}$ as given in [40]. Let $y$ be a representative of a $\bar{G}_{\sigma}$-class corresponding to this pair. Inspecting [40] tells us that $C_{\bar{G}}(y)^{0}$ is of type $A_{2}^{2}T_{2}$, and in particular, $|C_{\bar{G}_{\sigma}}(y)|=\Phi_{3}(q)|{\rm SL}_{3}(q)|^{2}$, where $\Phi_{3}(q)=q^{2}-q+1$ denotes the third cyclotomic polynomial. We now turn to [66] and we find that the entry corresponding to the class of $y$ is the one with label $(i,j,k)=(12,1,5)$, and from there we can read the number of classes.
@@
$\beta\leqslant\sum_{i=1}^{s}|x_{i}^{\bar{G}_{\sigma}}|\cdot F(x_{i},q)^{6}$
where $x_{1},\ldots,x_{s}$ represent the prime order unipotent classes of $G$. In this way, we find that
@@
We finally compute a bound on $\gamma$. Let $\gamma_{1}$ denote the contribution from involutory field and graph-field automorphisms, $\gamma_{2}$ denote the contribution from the odd order field automorphisms and $\gamma_{3}$ denote the contribution from graph automorphisms.
@@
where we recall that $\bar{H}(q)=\bar{H}_{\sigma}$ and $\bar{H}^{\epsilon}(q^{1/r})=C_{\bar{H}_{\sigma}}(x)$. Note that this implies that ${\rm fpr}(x,G/H)$ is maximised when $|\bar{H}(q)|$ (and consequently $|\bar{H}^{\epsilon}(q^{1/r})|$) is maximised, and we certainly have $|\bar{H}(q)|\leqslant|P_{1}(q)|$, where $P_{1}$ is a $P_{1}$-parabolic subgroup of $E_{6}$. Moreover, we note that $|{}^{2}E_{6}(q)|>|E_{6}(q)|$ for all $q\geqslant 3$. Therefore, if $q\geqslant 5$ we find
$\gamma_{1}\leqslant\frac{3+\epsilon}{2}\cdot\left(\frac{|E_{6}^{\epsilon}(q)|}{|{}^{2}E_{6}(q^{1/2})|}\right)^{-5}\left(\frac{|P_{1}(q)|}{|P_{1}(q^{1/2})|}\right)^{6}$
@@
Finally, we infer from Propositions 8.1, 8.2, and 9.5 that if $x$ is an involutory graph automorphism, then $f(x,q)\leqslant q^{-5}$ if $C_{\bar{G}}(x)=F_{4}$ and $f(x,q)\leqslant 12q^{-10}$ if $C_{\bar{G}}(x)$ is not of type $F_{4}$. If $q\geqslant 5$, then we find
@@
One then checks that
@@
Finally, if $q=3$ or $4$, then by only considering the contribution by involutory graph automorphisms (and field and graph-field automorphisms if $q=4$), we compute $\gamma\leqslant 0.025$ using Magma.